### v2

# Self-Training Elicits Concise Reasoning in Large Language Models

Abstract
Chain-of-thought (CoT) reasoning has enabled large language models (LLMs) to utilize additional computation through intermediate tokens to solve complex tasks. However, we posit that typical reasoning traces contain many redundant tokens, incurring extraneous inference costs. Upon examination of the output distribution of current LLMs, we find evidence on their latent ability to reason more concisely, relative to their default behavior. To elicit this capability, we propose simple fine-tuning methods which leverage self-generated concise reasoning paths obtained by best-of-N sampling and few-shot conditioning, in task-specific settings. Our combined method achieves a 30% reduction in output tokens on average, across five model families on GSM8K and MATH, while maintaining average accuracy. By exploiting the fundamental stochasticity and in-context learning capabilities of LLMs, our self-training approach robustly elicits concise reasoning on a wide range of models, including those with extensive post-training.
Self-Training Elicits Concise Reasoning in Large Language Models
Tergel Munkhbatthanks: Equal contribution. Corresponding author.    Namgyu $**{}^{\text{*}}$    Seo Hyun $**{}^{\text{*}}$    Yongjin Yang    Yujin Kim    Se-Young Yun KAIST AI {tergelmunkhbat, itsnamgyu, shkimsally, yunseyoung}@kaist.ac.kr https://github.com/TergelMunkhbat/concise-reasoning
1 Introduction
Chain-of-thought (CoT) reasoning has significantly improved the ability of large language models (LLMs) to perform complex tasks Wei et al. (2022b). The effectiveness of CoT reasoning has been attributed to the additional computation allocated during inference, as each intermediate reasoning token enables the model to perform an additional forward pass through its parameters Nye et al. (2021); Wei et al. (2022b). On the other hand, this inherently incurs additional inference cost and latency, roughly proportional to the number of output tokens Agrawal et al. (2024); Ho et al. (2024).
Figure 1: Potential of concise reasoning in current LLMs. Examples of reasoning samples with different lengths and the normalized length distribution of correct reasoning paths from Llama-3.1-8B on GSM8K. Typical samples from the original model distribution marked in blue, as shown in the blue box, contain redundant context as marked in gray. We fine-tune the model with relatively concise reasoning paths within the subregion of the model output distribution marked in orange. This elicits the model to reason more efficiently, as shown in the orange box, retaining only the essential parts while still leading to the correct answer.
We posit that current models often generate more tokens than necessary to accomplish the task, incurring extraneous inference costs. Typical reasoning chains sampled from the original model distribution, exemplified in the blue box of Figure 1, includes many tokens that do not contribute to the solution, specifically verbose explanations and repetitive phrasing. Similar observations have also been made in prior work Renze and Guven (2024); Zhang et al. (2024); Chiang and Lee (2024a).
We argue that the redundancy in reasoning chains of current models is not surprising. LLMs have not been explicitly trained to utilize intermediate tokens for reasoning in the most efficient manner. Rather, CoT reasoning is an emergent ability of LLMs Wei et al. (2022a) first uncovered through novel few-shot prompting Wei et al. (2022b). Post-hoc analysis suggests that reasoning ability is derived from procedural knowledge in pretraining data Ruis et al. (2024), which is not optimized for conciseness. Therefore, it is natural that the default behavior of LLMs does not optimize for efficient use of tokens for reasoning. However, this does not preclude the possibility that LLMs possess the ability to reason more efficiently.
Indeed, the distribution of reasoning path lengths shown in Figure 1 reveals a compelling insight: the presence of shorter, more efficient reasoning paths within the model’s output distribution, marked in orange, suggests a latent capacity for concise reasoning. This observation motivates our central hypothesis: that by selectively leveraging these existing concise examples, we can fine-tune the model to elicit this latent ability and shift its default output distribution towards more efficient reasoning.
In this paper, we propose a simple yet effective fine-tuning method to elicit efficient, concise reasoning in models for a given target task. We note that existing zero-shot prompting methods fails to reliably elicit concise reasoning, exhibiting near negligible impact on task-specialized models (Section 2.2). On the other hand, we can leverage best-of-N (BoN) sampling and few-shot conditioning (FS) to reliably generate concise training data (Section 3). We can then apply standard fine-tuning to distill the length-reducing benefits of BoN sampling and few-shot prompting back into the model itself while avoiding their associated overheads at inference time.
Across GSM8K and MATH datasets and a wide variety of model families, our unified few-shot conditioned best-of-N sampling (FS-BoN) method achieves a substantial reduction in output length of 30% on average, a 2.4x improvement over previous fine-tuning baselines De Sabbata et al. (2024), while preserving overall accuracy. Notably, our analysis shows that trained models adaptively adjust output length based on question complexity, preserving detail for difficult problems while simplifying responses to easier ones. We confirm that these results remain consistent across various model scales. These results strongly suggest that fine-tuning with carefully curated, self-generated data effectively can unlock latent concise reasoning abilities within LLMs, leading to significantly more cost-effective inference for complex tasks.
2 Preliminary Investigation
Avg. Relative Ll-3B Ge-2B Qw-3B QM-1.5B DM-7B
Method Acc Len Acc Len Acc Len Acc Len Acc Len Acc Len
Default Pang et al. (2024) 100.0 100.0 77.3 220.0 65.4 200.1 84.4 301.1 82.8 305.9 80.4 182.3
Be Concise Renze and Guven (2024) 99.9 88.5 78.2 168.4 64.2 163.1 83.8 258.2 82.9 304.1 80.7 180.6
Est. Budget Han et al. (2024) 79.8 121.4 73.2 243.6 54.5 124.7 22.2 29.6 80.2 609.9 78.5 409.3
Fixed Budget Nayab et al. (2024) 89.9 67.8 71.7 129.7 56.5 129.9 63.8 68.8 81.7 298.1 77.4 173.1
Hand Crafted 1 97.4 81.1 77.9 179.7 60.4 123.6 80.1 186.5 81.7 306.1 80.8 182.3
Hand Crafted 2 98.3 77.1 79.2 143.6 62.6 144.9 78.8 146.4 82.5 303.5 80.6 182.2
Hand Crafted 3 91.9 68.4 73.6 101.7 50.7 127.9 74.0 88.6 81.5 307.4 80.7 185.5
Hand Crafted 4 61.3 48.8 35.0 49.5 29.8 49.5 14.5 7.7 81.2 297.0 80.7 177.1
Table 1: Inconsistency of zero-shot prompting methods. Zero-shot prompting methods for concise reasoning on GSM8K. Relative accuracy (%) and length (%) compared to default prompting are shown, averaged over five main models: Llama-3.2-3B, Gemma-2-2B, Qwen2.5-3B, Qwen2.5-Math-1.5B, and DeepSeekMath-7B. Absolute accuracy (%) and length (tokens) are reported for each individual model. Values in red indicate significant relative accuracy drop greater than 5%. Values in gray indicate negligible relative length reduction less than 5%.
In this section, we probe the potential of current models to perform concise reasoning (Section 2.1) and investigate whether zero-shot prompting can reliably elicit such behavior (Section 2.2). We start with a definition of concise reasoning and key experimental setup.
Figure 2: Models already exhibit relatively concise reasoning in their output distributions to differing degrees. Distribution of output token lengths in reasoning paths of GSM8K, normalized by the average length of correct reasoning paths for each question, across five main models. Reasoning paths to the left of the gray line are relatively shorter than the model’s default output (Section 2)
Concise reasoning
We define concise reasoning in relative terms: for a given model, concise reasoning is to correctly reason through a given problem using less output tokens, compared to its default output. What is the default output? For our preliminary investigation, we measure the default output length of a model on a given question by taking the average length of correct paths over multiple stochastic generations. We use this measure instead of greedy decoding length, as the greedy path may be incorrect. We also consider the greedy output length as a noisy estimate, as it is highly sensitive to superficial input perturbations and numerical errors (Holtzman et al., 2019).
Key experimental setup
We analyze a wide range of moderately sized models, including both instruction-tuned language models and models specifically fine-tuned for mathematical reasoning, to assess robustness. We consider two reasoning tasks of moderate difficulty to assess preservation of model performance: GSM8K and MATH. We explain our full experimental setup in Section 4.
2.1 Reasoning Length Distribution
To assess the potential of current models to elicit concise reasoning, we analyze the extent of concise reasoning present within their output paths. Following our definition of concise reasoning in this context, we analyze the normalized length of correct reasoning paths, which is obtained by dividing the length (number of output tokens) of each path by the average length of correct paths for that question. We stochastically sample 16 paths per question on the GSM8K training dataset and visualize the distribution of normalized lengths using kernel density estimation (Scott, 2015). Alternative analysis on absolute token count and raw character count, as well as results on the MATH dataset, are present in Appendix Section D.1.
Models can reason with fewer tokens than their default to varying degrees.
In Figure 2, the distribution of normalized lengths show substantial mass below 100%, indicating frequent occurrences of solutions shorter than the default length. This suggests that models have the capability to produce more concise solutions than their average output would suggest. Notably, DeepSeekMath-7B, which already uses the least tokens in absolute terms (see Figure 8), produces correct solutions using less than half of its average token count in 8.37% of cases, demonstrating substantial potential for more concise reasoning. We find similar patterns on the MATH dataset, detailed in Section D.1.
2.2 Efficacy of Zero-Shot Prompting
While several zero-shot prompting methods have been proposed to elicit concise reasoning, evaluation has been limited to few models. We reassess their impact on output length and reasoning performance across diverse model families, evaluating accuracy and output length, relative to that of a strong default prompt (Pang et al., 2024). Be Concise (Renze and Guven, 2024) appends the phrase ‘be concise’ to the baseline prompt. Estimated Budget (Nayab et al., 2024) prompts the model to ‘use less than N tokens,’ where N is determined by the model itself. Fixed Budget (Han et al., 2024) prompts the model to ‘limit the answer length to 100 words.’ In addition, we introduce four novel prompts optimized on Llama-3.2-1B, as our best-effort attempt at prompt optimization. Full prompts and prompting details are outlined in Appendix F.
Output length reduction typically comes at the cost of accuracy.
As shown in Table 1, most zero-shot prompting methods reduce reasoning length but often lead to notable accuracy loss. For example, the Fixed Budget prompt shortens reasoning by 32.2% but lowers accuracy by 10.1% relative to the baseline, on average. Our hand-crafted prompts confirm that greater length reduction generally results in higher accuracy degradation.
Zero-shot prompting is less effective on task-specific models.
Zero-shot prompting methods show inconsistent performance across model families. While some prompts are effective on general-purpose models like Llama-3.2, they fail to elicit concise reasoning on math-specialized models like Qwen2.5-Math (Table 1). This suggests the internal representations of task-specific models are less responsive to zero-shot prompting. Similar trends are observed on the MATH dataset in Table 6.
3 Method
To address the limitations of zero-shot prompting, we propose a simple yet robust fine-tuning (FT) approach to reliably elicit concise reasoning in LLMs while preserving accuracy. We focus on self-training approaches, to unlock and refine the latent concise reasoning abilities of current LLMs, observed in Section 2.1. Self-training not only removes external dependencies, but we posit that it helps preserve reasoning capability, as the training data originates from the model’s own distribution.
3.1 Naive Best-of-N Sampling (BoN)
We first consider naive BoN sampling to collect relatively concise reasoning samples from the original output distribution, corresponding to the probability mass toward the left side of the distribution in Figure 2, for fine-tuning. Specifically, we generate $N$ reasoning paths for each question in the original training dataset and select the shortest correct reasoning path for each question [^1] Note that questions without any correct reasoning paths are excluded from the fine-tuning dataset.. In contrast to selecting the shortest subset of correct reasoning paths from a combined pool of questions, our question-wise selection scheme ensures supervision across a wide range of difficulty levels, as difficult questions may require longer absolute reasoning lengths. Corresponding experimental results are in Section D.3.
3.2 Few-Shot Conditioning For Effective Reduction
While naive BoN sampling is a straightforward approach, its sample inefficiency Xiang et al. (2025) makes it infeasible to achieve significant length reduction beyond a certain point. Figure 3 shows a log-linear relationship between $N$ and output length reduction, indicating that length reduction through BoN incurs exponential generation costs.
Figure 3: Sample efficiency of generation methods for self-training. Average relative length of reasoning paths collected by each method from 64 random questions in the training set. Relative length here is measured compared to the first path generated for naive BoN (only correct paths). Results are averaged over GSM8K and MATH across 4 models excluding DeepSeekMath-7B.
Few-shot conditioned sampling (FS)
To mitigate the sample inefficiency of naive BoN, we can leverage few-shot prompting to bootstrap the reduction in output length. We consider three sources for few-shot exemplars: human-annotation (FS-Human), proprietary frontier LLMs (FS-GPT4o), and self-generated samples (FS-Self). We use human annotated examples from Wei et al. (2022b), as they are readily available and demonstrate very concise reasoning. We provide additional details on few-shot example acquisition in Appendix B.
We find that few-shot prompting reliably elicits concise generations across all models considered in our study. As few-shot learning is a fundamental capability of LLMs (Brown et al., 2020), we expect this to further generalize to other models. Horizontal lines in Figure 3 shows that sampling with 8-shot conditioning reduces the length of reasoning paths significantly. Notably, FS-Human elicits more reduction than that achieved by BoN sampling with $N=256$.
Few-shot conditioned BoN sampling (FS-BoN)
We can maximize length reduction by jointly applying BoN sampling on top of few-shot conditioning. We consider GPT-4o generated examples for this approach (FS-GPT4o-BoN), based on their strong accuracy preservation and length reduction in standalone few-shot conditioned self-training (see FS-GPT4o in Table 2). Figure 3 shows that the improvements from few-shot conditioning and BoN sampling are largely independent and additive, resulting in significant length reduction.
Why do we need (self-)training?
Directly applying BoN and few-shot prompting at test time, without self-training, incurs prohibitive inference overhead due to repeated sampling and long prompts (Table 5). This defeats our ultimate goal of mitigating the inference costs of reasoning. On the other hand, self-training internalizes the length reduction benefits, allowing the model to reason concisely without additional inference-time overhead.
3.3 Sample Augmentation for Accuracy Boost
While few-shot prompting elicits concise reasoning, its adaptability is limited due to the small number of given examples. It may (1) prohibit generation of correct paths for very complex questions that require longer reasoning paths, while (2) eliciting extraneous steps that are not required for very easy questions. To address this, for each question, we augment the set of {1, $N$} sample(s) generated for FS and FS-BoN, respectively, with $N$ samples generated for naive BoN, and select the shortest correct path from the combined set. We find that this retains the length reduction from FS and FS-BoN while better preserving accuracy, likely due to better coverage of difficult questions. We apply this augmentation by default to few-shot conditioned methods, with its effectiveness ablated in Figure 4.
4 Experimental setup
Models
To account for realistic task-specific deployment settings, we select recent moderately sized post-trained models. We also consider math-specialized models to evaluate on models that have been optimized for specific task domains. We consider five main models for our key experiments: Llama-3.2-3B Dubey et al. (2024), Gemma-2-2B Team et al. (2024), Qwen2.5-3B Yang et al. (2024a), Qwen2.5-Math-1.5B Yang et al. (2024b), and DeepSeekMath-7B Shao et al. (2024). We investigate scaling on Llama-3.2-{1B, 3B} and Llama-3.1-8B.
Tasks and datasets
We focus on challenging reasoning tasks where (1) CoT reasoning significantly improves model performance, (2) only the final answer is relevant, and (3) models achieve moderate performance. Reasoning length reduction is desirable under the first and second conditions as it can reduce inference latency without affecting utility. The third condition is necessary to assess accuracy preservation. We consider two mathematical reasoning datasets: GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), where the models achieve accuracies of 40–90% and 20–70%, respectively. For evaluation, we utilize the test set of the GSM8K and MATH500 dataset (Lightman et al., 2024). We explain details in Section C.1.
Evaluation metrics
We evaluate methods using two primary metrics: accuracy and length. Accuracy is evaluated using Python-based parsing code, described in Section C.2. Length is defined as the average number of output tokens in all reasoning paths, including incorrect ones, as output tokens incur inference costs in deployment scenarios regardless of their correctness. We focus on number of output tokens for clear comparison, as number of input tokens are similar across methods, and output tokens affect wall-clock latency to a higher degree Agrawal et al. (2024). We further justify this choice in Section C.3. We also employ relative accuracy and relative length metrics to better evaluate how well each method elicits concise reasoning while maintaining accuracy. Specifically, relative accuracy is the ratio of the given method’s accuracy to the baseline accuracy, and relative length is the ratio of the given method’s average length to the baseline average length, using a strong baseline zero-shot prompt Pang et al. (2024). We use greedy decoding throughout evaluation to ensure reproducibility222 Note that this contrasts with Section 2 where we took the average length of correct samples over stochastic generations per question, for precise question-level length normalization. .
Baseline methods
For baselines, we consider zero-shot prompting (Section 2.2) and fine-tuning directly on ground-truth answers, and with external supervision from human- and GPT-4o-generated concise reasoning paths. For existing fine-tuning methods on concise reasoning, we reproduce Rational Metareasoning (RM) De Sabbata et al. (2024), which follows a similar approach to our naive BoN method but with two key differences: (1) a reward function that balances efficiency and utility (i.e., output length and accuracy), and (2) iterative training via expert iteration Zelikman et al. (2022).
Self-training budget allocation
The cost of self-training methods can be broken down into (1) generation and (2) fine-tuning. We aim to match the budget across methods in terms of generation (number of paths generated per question) since fine-tuning costs are relatively negligible (see Table 5). For naive BoN, we generate 16 paths per question. For RM De Sabbata et al. (2024), we generate $N=4$ paths per iteration across 4 iterations. We generate a single path for standalone few-shot conditioning (FS) and 16 paths for FS-BoN, both augmented with 16 additional paths from the default output distribution (used for naive BoN). Considering the large budget of FS-BoN with augmentation, we evaluate a Budget-Matched setting with 8 paths each from the few-shot conditioned distribution and default distribution, as shown in Table 2. Details are provided in Section C.4.
Dataset GSM8K MATH
Method Acc Len Rel. Acc Rel. Len Acc Len Rel. Acc Rel. Len
Zero-Shot Prompting
      Baseline 78.06 241.87 100.00 (0.00) 100.00 (0.00) 46.40 480.37 100.00 (0.00) 100.00 (0.00)
      Be Concise 77.98 214.87 99.85 (1.18) 88.46 (10.37) 47.76 446.09 102.71 (7.59) 92.66 (7.46)
      Hand Crafted 2 (ours) 76.72 184.13 98.27 (3.67) 77.10 (22.27) 46.84 404.85 101.62 (4.79) 85.26 (15.97)
FT - External Data
      Direct Answer 19.70 3.17 24.88 (5.03) 1.36 (0.40) 15.08 6.98 35.16 (10.34) 1.44 (0.73)
      Human CoT 65.73 127.85 83.82 (7.28) 54.95 (13.17) 33.88 243.54 75.61 (13.56) 53.14 (13.87)
      GPT4o CoT 76.36 156.24 97.65 (3.63) 67.60 (16.70) 40.44 399.80 90.52 (15.07) 87.21 (22.22)
FT - Best-of-N Self-Generation
      Naive BoN 77.12 214.22 98.79 (1.64) 87.17 (8.79) 47.64 433.26 101.74 (7.04) 89.89 (3.99)
      Rational Metareasoning 76.15 207.49 97.21 (5.74) 84.93 (5.09) 47.56 432.56 103.02 (6.56) 90.56 (5.25)
FT - Few-Shot Conditioned Self-Generation (ours)
      FS-Human 76.66 161.72 98.06 (3.28) 67.96 (16.62) 46.44 421.54 99.69 (6.97) 87.78 (5.98)
      FS-GPT4o 78.07 175.54 99.94 (1.69) 73.15 (13.49) 47.36 421.21 101.87 (5.33) 87.58 (6.60)
      FS-Self 77.27 190.03 98.86 (2.51) 77.51 (9.18) 48.00 426.67 102.67 (5.24) 88.50 (4.49)
FT - Few-Shot Conditioned Best-of-N Self-Generation (ours)
      FS-GPT4o-BoN 75.88 153.38 97.00 (4.11) 64.25 (16.66) 47.36 364.33 102.56 (6.24) 76.30 (10.56)
[…]
Table 2: Evaluation of zero-shot prompting, fine-tuning (FT), and our proposed methods on GSM8K and MATH datasets, averaged across five main models. We report absolute accuracy (%) and length (tokens), along with relative accuracy (%) and length (%) compared to baseline prompting. Values in parentheses show standard deviations.
5 Results
Question
A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?
Reasoning Paths
[Zero-Shot] To find the total number of bolts needed, we need to calculate the amount of white fiber first, since it’s half the amount of blue fiber. Step 1: Determine the amount of blue fiber needed. The robe takes 2 bolts of blue fiber. Step 2: Calculate the amount of white fiber needed. The robe takes half the amount of white fiber as blue fiber. Half of 2 bolts is $/2$ 2 = 1 bolt. Step 3: Add the amount of blue fiber and white fiber to find the total. Total bolts = blue fiber + white fiber. Total bolts = 2 bolts + 1 bolt. Total bolts = 3 bolts. The answer is 3.
[FS-GPT4o-BoN] The robe takes 2 bolts of blue fiber. It takes half that much white fiber, which is $/2$ 2 = 1 bolt. Add the blue and white fiber together: 2 + 1 = 3 bolts. The answer is 3.
Table 3: Example rationales generated from the original Llama-3.1-8B model (Zero-Shot) and the model fine-tuned using our method (FS-GPT4o-BoN). The question is randomly sampled from GSM8K. For the same question, one can observe that our rationale is much more concise, removing redundant parts that explain the question.
5.1 Main results
Our main results, presented in Table 2 and Figure 4, demonstrate the performance of our self-training methods against baseline approaches.
Naive BoN fine-tuning is effective but sample inefficient.
Naive BoN fine-tuning effectively reduces output length without significantly degrading model performance. This also holds true for Qwen2.5-Math-1.5B and DeepSeekMath-7B (Table 8 and Table 9), which failed to achieve length reduction through zero-shot prompting. However, the length reduction from naive BoN with $N=16$ is limited to 12% on average. Furthermore, as illustrated in Figure 3, achieving more compression with BoN becomes progressively less efficient.
Iterative baseline yields similar results as naive BoN fine-tuning.
Rational Metareasoning, an iterative baseline, yields similar relative length reduction and relative accuracy to BoN fine-tuning. This suggests that the utility reward proposed by De Sabbata et al. (2024) may not effectively achieve both accuracy gains and token length reduction.
Figure 4: Tradeoff between relative accuracy and length reduction for main methods. Results are averaged over GSM8K and MATH across five main models. Matching colors and shapes indicate the same FS prompt. FS conditioning without augmentation (†) are marked with lighter colors. Relative length reduction refers to 100 - relative length (%).
Few-shot conditioning outperforms BoN in length reduction.
The results demonstrate that few-shot conditioning achieves a greater relative length reduction compared to naive BoN, including math-specialized models (Table 8 and Table 9). This is in line with the superior length reduction of few-shot conditioning, compared to naive BoN as shown in Figure 3. Notably, self-training on generations conditioned on human-annotated examples (FS-Human) achieves an average relative length of 67.96% on GSM8K, compared to 87.17% with naive BoN.
Self-training better preserves accuracy than training with external data.
Table 2 shows fine-tuning with external data (FT-External Data) leads to a significant reduction in relative length but causes a severe drop in relative accuracy. Figure 4 further highlights the accuracy preservation of self-training: fine-tuning with external concise reasoning supervision from GPT-4o (FT-GPT4o) lies below the Pareto-curve of relative accuracy and relative length reduction, established by our self-training methods.
Few-shot conditioned BoN achieves best length reduction while maintaining accuracy.
FS-BoN elicits the largest length reduction among our self-training methods, while maintaining relative accuracy, on average. Notably, for math-specialized models, FS-GPT4o-BoN achieves the greatest reduction among all methods, except those fine-tuned on external data which greatly sacrifice the accuracy (Table 8 and Table 9).
Augmentation boosts accuracy for few-shot conditioning.
Figure 4 compares few-shot conditioning, i.e., FS and FS-BoN, with and without augmentation (†). Augmentation improves accuracy by providing solutions for previously unsolvable hard questions as discussed in Section 3.3. While augmentation may slightly affect reduction rates, they remain superior to naive BoN and RM. Even when matching the budget (Budget-Matched) with other self-training methods in Table 2, it achieves the greatest length reduction among them with minimal accuracy degradation. The effect of augmentation on training data length is analyzed in Section D.5.
Figure 5: Tokens are reduced adaptively according to question difficulty. Token reduction rate for each difficulty level on MATH, for 4 models individually and averaged. Relative length reduction refers to 100 - relative length (%).
5.2 Analysis
In this section, we analyze the length reduction effects in depth. We exclude DeepSeekMath-7B from quantiative analysis due to cost.
Tokens are reduced adaptively according to question complexity.
The MATH dataset’s difficulty levels range from 1 (basic algebra) to 5 (advanced calculus and complex mathematical reasoning). As shown in Figure 5, our method adaptively reduces tokens based on question difficulty, with higher difficulty leading to less reduction. The higher reduction (20%–40%) at easier difficulty levels (1–2) suggests that the original model outputs for these easier questions contained unnecessary tokens. This reveals a gap in current models’ ability to tailor their inference budget to problem complexity. Our method effectively closes this gap by reducing redundancy, allowing for more precise token allocation based on question difficulty.
Figure 6: Scaling study on baseline and few-shot conditioned self-training methods. Results are averaged over GSM8K and MATH for Llama 1B, 3B, and 8B. Relative length reduction refers to 100 - relative length (%).
Self-training maintains consistency across model scales.
We conduct a scaling study on Llama-3.2-1B, 3B, and Llama-3.1-8B to examine consistency across different model sizes (Figure 6). Overall, token reduction increases as the model size increases, while the maintenance of accuracy does not show a strong correlation with model size. RM exhibits lower reduction rates compared to our few-shot conditioned self-training methods across all models and shows a decrease in accuracy with increasing model size. Our standalone few-shot conditioning method (FS-GPT4o) also shows a similar trend in length reduction, but better preserves accuracy. Our joint FS-GPT4o-BoN method achieves the greatest reduction across all models, maintaining relative accuracy across different model sizes, especially in the largest 8B model.
Sample study
Table 3 presents qualitative examples of reasoning paths generated by the model before and after fine-tuning with our method. The original reasoning exhibits verbosity, containing redundant processes such as question confirmation and repeated instructions. In contrast, the reasoning generated by our method includes only the necessary steps, significantly reducing the number of tokens while still arriving at the correct answer. More examples are provided in the Appendix E.
Figure 7: Fine-tuning effectively transfers the length reduction to the model. Correlation between the relative length of train data and test output averaged over GSM8K and MATH across 4 models. Training length includes only correct solutions. Solid points represent test lengths including all generated outputs, while lighter points indicate test lengths of correct solutions only.
Length reduction is transferred through fine-tuning.
As shown in Figure 7, fine-tuning with shorter rationales results in shorter model outputs, showing a strong correlation between test and training lengths. We note that the length of test outputs (incorrect and correct) are longer than the length of training samples (only correct) on average. This is mainly because incorrect paths are generally longer than correct ones. We find a closer correspondence between train length and test length of correct samples only, indicated by the lighter datapoints. This discrepancy suggests the need to terminate incorrect paths early to minimize redundant inference overhead. We consider this for future work.
6 Discussion
Default reasoning behavior of LLMs
Previous research has shown that the CoT reasoning ability of LLMs originates from procedural knowledge in pretraining data Ruis et al. (2024). Modern LLM training pipelines utilize high-quality math, code, and synthetic reasoning data to enhance reasoning, but these do not promote conciseness Dubey et al. (2024); Yang et al. (2024a). Furthermore, recent ‘thinking’ models are reinforced to use additional tokens to improve reasoning performance, rather than save on token budget OpenAI (2024). Thus, current LLMs naturally exhibit redundant reasoning. We believe that incorporating concise reasoning supervision or rewards in training pipelines can be beneficial for model efficiency, especially for ‘thinking’ models with lengthy internal reasoning.
Lightweight fine-tuning for concise reasoning elicitation
In this paper, we focused on standard fine-tuning based on self-generated samples. Our analysis shows that LLMs already possess the ability to correctly reason in a relatively concise manner. Therefore, we posit that lightweight fine-tuning is sufficient to achieve significant reasoning length reduction, echoing the Superficial Alignment Hypothesis in the field of LLM alignment Zhou et al. (2023), Indeed, our post-hoc analysis in Figure 7 shows that standard fine-tuning can reduce the output length of models in proportion to that of training samples.
7 Conclusion
We tackle redundancy in CoT reasoning, hypothesizing LLMs possess a latent capacity for concise reasoning, evidenced by shorter correct reasoning paths. We introduce fine-tuning methods, leveraging self-generated data from BoN sampling and few-shot conditioning, to elicit this capacity. Our FS-BoN method significantly reduces reasoning length by 30% reduction while maintaining accuracy. This implies fine-tuning with curated self-generated data can reliably unlock latent concise reasoning, enabling more efficient inference.
Limitations and Future Work
Advanced Training Schemes
While our focus was on standard fine-tuning with self-generated samples, exploring advanced reinforcement learning (RL) based training schemes could potentially maximize efficiency further. Our preliminary experiments with expert iteration and BoN sampling (without a utility reward) showed promising results, boosting both accuracy and length reduction compared to non-iterative BoN. Further investigation of iterative or more advanced RL methods is warranted.
Few-Shot Prompting Exploration
Although we leveraged few-shot prompting to bootstrap length reduction during data generation, we did not explore advanced few-shot prompting methods as a middle ground between zero-shot and fine-tuning. While direct few-shot prompting with our prompts achieved comparable length reduction to fine-tuning on single-path few-shot generations, it incurred a slight accuracy loss. Critically, fine-tuning enabled us to incorporate BoN sampling and sample augmentation, leading to significantly greater length reduction without sacrificing accuracy. However, future work on integrating advanced exemplar selection Fu et al. (2022) and many-shot prompting Agarwal et al. (2024) could further enhance our fine-tuning approach.
Extended Scaling Studies
Our scaling study was limited to Llama 3.x models at 1B, 3B, and 8B parameters. While these model sizes are relevant for our task-specific setting, further empirical studies are needed to evaluate the effectiveness of our method on models exceeding 8B parameters. Investigating scaling trends across a wider range of model sizes is crucial for understanding the full potential of our approach.
Concise Reasoning in General LLMs
This study focused on task-specific fine-tuning. While existing zero-shot and fine-tuning methods often struggle with reliability and efficacy even in such settings, and our method proved effective, generalizing our approach to a broader range of tasks is an important direction. Exploring techniques like multi-task training could enable efficient reasoning without task-specific tuning. This could be particularly beneficial for ‘thinking’ models (Guo et al., 2025) where reasoning enhances the final response, similar to our math reasoning tasks but in a more general context. Furthermore, while prior work has explored reducing the number of reasoning stages Chen et al. (2024), our method focuses on reducing verbal redundancy within sentences, offering a complementary approach.
Acknowledgements
We would like to express our gratitude to Taehyeon Kim, Murad Aghazada, Sunwoo Hong, and Minu Kim for their valuable feedback and assistance throughout this research project. Their insightful comments and suggestions significantly contributed to the quality of this work.
Appendix A Related work
A.1 Chain-of-thought Reasoning
CoT reasoning enables LMs to solve diverse tasks (Wei et al., 2022b; Kojima et al., 2022) by leveraging an arbitrarily large amount of computation for next-word prediction (Merrill and Sabharwal, 2023; Nowak et al., 2024; Xiang et al., 2025). Recently, additional lines of research have expanded (OpenAI, 2024; Guo et al., 2025) basic CoT by adding in-context search (Ye et al., 2024; Kumar et al., 2024), often termed as meta-CoT (Xiang et al., 2025) or long-CoT (Yeo et al., 2025). This paper focuses on original CoT framework to investigate the simplest ways to reduce the number of tokens, which could potentially be applied to long-CoT as well.
A.2 Concise Reasoning
Many studies have shown that LLMs tend to overthink during CoT, often including unnecessary words in their reasoning (Zhang et al., 2024; Chiang and Lee, 2024b; Wang et al., 2024), which can negatively impact efficiency. However, simply reducing the number of tokens often leads to significantly lower performance  (Merrill and Sabharwal, 2023; Jin et al., 2024). To address this, some works have explored prompting techniques to elicit concise reasoning (Renze and Guven, 2024; Nayab et al., 2024; Han et al., 2024), but these methods are highly model-dependent. Another line of work has focused on fine-tuning methods to balance efficiency and reasoning quality (De Sabbata et al., 2024), but these approaches are insufficient in reducing token usage. In this work, we investigate how to efficiently generate concise reasoning samples and use them for fine-tuning, improving model efficiency without sacrificing accuracy.
Several concurrent works also address redundancies in LLM outputs. Some approaches Chen et al. (2024); Luo et al. (2025) focus on how o1-like models overthink and propose methods for pruning extraneous reasoning steps. Another approach, TokenSkip Xia et al. (2025), first identifies and discards low-importance tokens from self-generated CoT outputs while preserving crucial ones, then fine-tunes the model on these compressed examples to enable efficient inference. In contrast, our method elicits concise reasoning more simply yet effectively by training on a self-generated dataset of shorter rationales, using best-of-N sampling and few-shot conditioning.
Appendix B Methodological Details
B.1 Self-Generation of Few-Shot Exemplars (FS-Self)
We devise a systematic approach to obtain few-shot exemplars using the default prompt (Appendix F.1) of concise reasoning from the target model itself, motivated by BoN sampling, through a two-phase process: sampling and selection.
Sampling Phase
In the sampling phase, 128 questions were randomly selected from the training set, and for each selected question, 128 diverse reasoning paths were generated with a sampling temperature of $T=0.7$. This process created a diverse set of candidate paths for both the GSM8K and MATH datasets.
Selection Phase
The selection process was executed through the following steps:
  1. 1.
    Sorting candidate paths by token count in ascending order.
  2. 2.
    Filtering only correct samples by comparing oracle labels with the models’ answers using the parsing code (Section C.2).
  3. 3.
    Sequential validation using GPT-4o with the following specialized prompt was performed until eight valid and concise examples were obtained. To ensure question uniqueness, alternative solutions from previously selected questions were excluded.
You are a math reasoning validator. Given a math problem, its answer, and a proposed rationale: First, determine if the problem requires step-by-step reasoning to solve.
Then, validate if the rationale contains explicit step-by-step reasoning that correctly solves the problem.
Question: [question here]
Expected Answer: [label here]
Proposed Rationale: [rationale here]
Respond with ONLY ACCEPT or REJECT:
ACCEPT - if the problem requires step-by-step reasoning AND the rationale contains explicit step-by-step reasoning that correctly leads to the answer.
REJECT - if the problem is too simple (can be solved in one step).
REJECT - if the rationale lacks explicit reasoning steps or contains incorrect logic.
REJECT - if the rationale only states the answer.
Dataset-Specific Selection Criteria
For GSM8K, eight valid examples were selected using the above selection phase. An example is shown in Table 17. For MATH, to avoid selecting only very easy questions, the selection was distributed across different categories: two examples from algebra (the largest category) and one example from each of the other categories (geometry, intermediate algebra, prealgebra, number theory, counting and probability, and precalculus). An example is shown in Table 18.
B.2 Generation of Few-Shot Exemplars From GPT-4o (FS-GPT4o)
To obtain few-shot exemplars of concise reasoning from GPT-4o, we use our zero-shot prompt Hand Crafted 3 (Appendix F.1) to generate short reasoning paths. For GSM8K, we randomly sample eight questions from the training set. For MATH, we randomly select two questions from the largest category, Algebra and one from each of the other categories (Geometry, Intermediate Algebra, Prealgebra, Number Theory, Counting & Probability, Precalculus), all from the training set. We manually filter low-quality outputs to ensure the quality of few-shot examples. However, we found that GPT-4o reliably followed the zero-shot prompt in the majority of cases. Generated examples of GPT-4o are in Tables 15 and 16.
Appendix C Experimental Details
C.1 Datasets
A summary of the datasets used in our experiments, along with their original licenses, is provided in Table 4. Both datasets are in English and focus on mathematical reasoning tasks. GSM8K contains grade school math word problems requiring multi-step reasoning, while MATH covers more advanced problems across various categories (algebra, geometry, precalculus, etc.) with different difficulty levels. Both datasets use straightforward language and standard mathematical notation. The original train/test splits are used for fine-tuning and evaluation. Results for MATH are reported on the MATH-500 subset, following the previous work Lightman et al. (2024).
Dataset Train Test License Reference
GSM8K 7,473 1,319 MIT Cobbe et al. (2021)
MATH 7,500 500 MIT Hendrycks et al. (2021)
Table 4: Dataset statistics for GSM8K and MATH.
C.2 Answer Parsing
For accuracy, we extract and verify the final numerical answer using Python-based parsing code. We build previous implementations for GSM8K (Kojima et al., 2022) and MATH (Lai et al., 2024), with additional rules to ensure consistent parsing across all models. Our implementation additionally handles mathematical expressions by standardizing numerical formats, removing units and mathematical notation, and cleaning model-specific formatting to ensure consistent evaluation across models.
During our experiments, we observed that different models require different parsing approaches to extract answers correctly for the MATH dataset. We implemented a model-specific parsing strategy where Gemma-2-2b and Qwen2.5-3B models utilize an enhanced parser, while all other models (including Llama, DeepSeekMath, and Qwen2.5-Math) use our standard parser. This adjustment led to more accurate model performance evaluation. All experiments use the appropriate parser for each model to ensure consistent evaluation.
C.3 Justification For Length Metric
We define our length metric for model evaluation as the average number of output tokens in model responses across the entire evaluation set. For clear comparison, we do not include input tokens because (1) the number of input tokens do not differ significantly between zero-shot prompting and direct prompting on fine-tuned models, but more importantly (2) the real-world wall-clock latency incurred by each output token is significantly greater than that of each input token. The greater latency of output tokens stems from the sequential nature of their processing (the decode stage). Input tokens, in contrast, are processed in parallel (the prefill stage). The decode stage is heavily memory-bound, requiring parameter fetching for every forward pass, which significantly limits compute utilization Agrawal et al. (2024). This difference in processing is also reflected in the pricing structures of commercial LLM API providers, where input tokens are cheaper than output tokens [^3] https://openai.com/api/pricing/.
C.4 Generation and Fine-Tuning
We implemented all model fine-tuning using the HuggingFace Trainer library and conducted generation using vLLM for optimized inference. For evaluation, we use greedy decoding, while for distribution analysis and training data generation we use temperature sampling with $T$=0.7, following Cobbe et al. (2021); Wang et al. (2023). We generate up to 512 output tokens on the GSM8K dataset and up to 1024 output tokens on the MATH dataset, following the token limits established in prior works Dubey et al. (2024); Ren et al. (2024). For fine-tuning, we use a batch size of 16 and train for one epoch with a learning rate of 1e-5. The fine-tuning phase requires at maximum 469 training steps, making it computationally modest compared to the generation phase, as demonstrated in Table 5. For comparison with the iterative fine-tuning baseline MR De Sabbata et al. (2024), we limit to 4 iterations as the authors observe performance generally plateaus after iteration 4. All experiments are conducted using bfloat16 precision for both generation and fine-tuning. We conduct our experiments on 8 NVIDIA H100 GPUs. Our total usage for main experiments is approximately 1,000 H100 GPU hours.
Appendix D Additional results
D.1 Full Analysis on Reasoning Length Distribution (Section 2.1)
We analyze the distribution of output lengths across different model families using three metrics. First, we examine the absolute output token count distribution, which directly reflects the computational cost during inference. Second, we analyze the raw character count distribution to verify our findings are consistent across different tokenization schemes. Third, we study the normalized output token count, where each solution’s output length is divided by the mean output length of correct solutions for that specific question, allowing us to identify relatively concise solutions.
Figure 8 reveals significant variations in how different model families approach mathematical reasoning. On GSM8K, DeepSeekMath-7B demonstrates notably concise reasoning with a mean output length of 179 tokens, while Qwen2.5-3B uses substantially more tokens, averaging 298 tokens per solution. This pattern persists in the MATH dataset, though with generally longer output lengths due to the increased complexity of the problems.
Approach Time
Default zero-shot generation (16 paths) 1h 0m 38s
GPT4o few-shot generation (16 paths) 1h 29m 54s
Model training 2m 24s
Table 5: Generation and training time comparison on a single H100 GPU for GSM8K using Llama-3.2-3B.
(a) GSM8K
(b) MATH
Figure 8: Distribution of absolute output token counts across model families on GSM8K and MATH datasets. The distributions show significant variation in output token usage patterns between different models, with some models consistently using more tokens than others for correct solutions.
The presence of a small peak at 512 tokens in some distributions is an artifact of the maximum output length limit used during generation, following standard practices in the field.
Figure 9 confirms that our observations about output token usage patterns are not merely artifacts of different tokenization schemes. The relative patterns between models remain consistent, suggesting that the observed differences in output length are fundamental to how each model approaches problem-solving, rather than being an artifact of tokenization choices.
The character-level distributions also provide additional insight into the actual verbosity of solutions, as they measure the raw text length without the influence of tokenization strategies.
(a) GSM8K
(b) MATH
Figure 9: Distribution of character counts across model families. The character-level analysis confirms that the observed patterns in output token usage are consistent when measured in raw UTF-8 characters, suggesting that the findings are not artifacts of different tokenization schemes.
(a) GSM8K
(b) MATH
Figure 10: Distribution of normalized output token counts, where each solution’s output length is divided by the mean output length of correct solutions for that question. The distributions reveal that models can produce valid solutions using significantly fewer tokens than their typical output length, though this capability varies across model families.
Avg. Relative Ll-3B Ge-2B Qw-3B QM-1.5B DM-7B
Method Acc Len Acc Len Acc Len Acc Len Acc Len Acc Len
Default Pang et al. (2024) 100.0 100.0 44.4 511.8 24.0 376.0 56.8 605.5 69.8 551.4 37.0 357.3
Be Concise Renze and Guven (2024) 102.7 92.7 47.2 477.5 22.2 307.1 60.2 541.9 68.0 550.2 41.2 353.8
Est. Budget Han et al. (2024) 80.7 132.7 45.6 907.4 16.2 255.1 20.4 125.7 70.0 1070.0 36.0 726.3
Fixed Budget Nayab et al. (2024) 88.1 70.8 42.6 413.1 22.0 193.6 33.0 170.3 69.6 536.7 35.2 344.8
Hand Crafted 1 (ours) 100.5 86.6 46.8 481.1 20.2 241.6 56.8 446.3 68.0 548.8 42.8 362.3
Hand Crafted 2 (ours) 101.6 85.3 46.6 445.9 24.0 275.7 54.8 393.8 68.8 543.7 40.0 365.2
Hand Crafted 3 (ours) 84.9 80.1 24.4 283.0 22.0 365.7 46.0 272.9 67.6 552.3 37.0 367.7
Hand Crafted 4 (ours) 63.9 45.9 13.2 65.7 13.0 88.7 16.2 9.4 69.4 535.5 39.8 337.9
Table 6: Inconsistency of zero-shot prompting methods on MATH. Zero-shot prompting methods for concise reasoning on MATH. Relative accuracy (%) and length (%) compared to default prompting are shown, averaged over five main models: Llama-3.2-3B, Gemma-2-2B, Qwen2.5-3B, Qwen2.5-Math-1.5B, and DeepSeekMath-7B. Absolute accuracy (%) and length (tokens) are reported for each individual model. Values in red indicate significant relative accuracy drop greater than 5%. Values in gray indicate negligible relative length reduction less than 5%.
Figure 10 shows a particularly interesting phenomenon. All models demonstrate the ability to solve problems with significantly fewer output tokens than their typical output length, though this capability varies across model families. The normalization process, which divides each solution’s output length by the mean output length of correct solutions for that specific question, reveals that some models are more flexible in producing concise solutions than others.
The presence of significant probability mass below 100% in the normalized distributions highlights an intriguing aspect of model behavior: while models may tend toward verbose outputs, they are not constrained to this verbosity and can successfully solve problems with more compact reasoning in some cases.
D.2 Full Analysis on Efficacy of Zero-Shot Prompting Methods (Section 2.2)
Table 6 shows evaluation results for zero-shot prompting methods on MATH. Prompting methods exhibit similar limitations as those on the GSM8K dataset.
D.3 Impact of Question-Wise Selection
We compare the performance of FS-GPT4o-BoN when selecting the shortest samples per question versus selecting the shortest samples across all questions. Both methods are tested without augmentation. The average relative accuracy across all models and datasets is 99.03% with question-wise selection and 96.03% without it, while the average relative length is 70.54% and 64.33%, respectively. Although selecting the shortest samples across all questions reduces relative length, it comes at the cost of accuracy. Notably, the accuracy of Gemma-2-2B drops significantly from 90.53% with question-wise selection to 75.81% without it. Therefore, we adopt question-wise selection to maintain question diversity and accuracy.
D.4 Full Result Tables on Main Methods
We present full results on the accuracy and reasoning length elicited by main methods for GSM8K in Table 8 and MATH in Table 9. We show absolute accuracy and output token length for individual models and show average over relative accuracy and relative length for aggregates across models.
D.5 Impact of Augmentation on Training Data Length
Table 7 demonstrates the impact of augmentation on training data length. Augmentation effectively expands the dataset by leveraging samples from zero-shot BoN. This is achieved by incorporating previously unsolvable long solutions, while also replacing solutions that can be solved with shorter solutions with zero-shot prompting. Consequently, the average solution length remains relatively stable, while the number of correct samples for fine-tuning increases, leading to improved model accuracy.
Appendix E Sample Study
Tables 10-13 present additional examples of generated reasoning paths from both the initial model and the model fine-tuned using FS-GPT4o-BoN. We primarily showcase results for the general model Llama-3.2-3b and the math-fine-tuned model Deepseek-7b, evaluated on the GSM8k and MATH500 datasets. One can observe a consistent reduction in token length in our fine-tuned models, achieved through decreased verbosity. This demonstrates the effectiveness of our method in improving inference efficiency.
Appendix F Prompts
F.1 Zero-Shot Prompts
For each question, we use the following instructions as the system input to the language model if the model’s chat template includes a system role. If no system role exists, we append the prompt to the user input instead.
F.1.1 Prompting Template
We used the default chat template for all models except Llama. For Llama, we modified its default chat template by removing the automatically added date information:
Cutting Knowledge Date: December 2023
Today Date: 23 July 2024
This modification was made to ensure fair comparison across all models by maintaining consistent system prompt formatting.
F.1.2 System Inputs:
Default Pang et al. (2024)
Your task is to answer the question below. Give step by step reasoning before you answer, and when you’re ready to answer, please use the format ‘The answer is’.
Be Concise Renze and Guven (2024)
Your task is to answer the question below. Give step by step reasoning before you answer, and when you’re ready to answer, please use the format ‘The answer is’. Be concise.
Est. Budget Han et al. (2024)
Budget Estimation: Analyze the given question and estimate the minimum number of tokens required to generate a complete and accurate response. Please give the response by strictly following this format: [[budget]], for example, Budget: [[12]].
Token-budget-aware CoT: Let’s think step by step and use less than {budget_estimate} tokens.
Fixed Budget Nayab et al. (2024)
Let’s think a bit step by step and limit the answer length to 100 words.
Hand Crafted 1 (ours)
Your task is to answer the question below. Give step by step reasoning before you answer, and when you’re ready to answer, please use the format ‘The answer is’. You don’t need unnecessary explanations; you can solve problems using only essential words and expressions. Summarize your thought process as simply as possible and provide your answer. Do not generate only the final answer.
Hand Crafted 2 (ours)
Your task is to answer the question below. Give step by step reasoning before you answer, and when you’re ready to answer, please use the format ‘The answer is’. Use less words and more compact expressions to be concise.
Hand Crafted 3 (ours)
Give concise step by step reasoning before you answer. Only retain key steps such as names, objects, numbers and mathematical operations. Use short plain words. Don’t use any formatting such as emphasis, lists, or enumeration. Make sure that the intermediate results are presented in the order that they are calculated. After your reasoning, please use the format ‘The answer is’ to answer.
Hand Crafted 4 (ours)
Your task is to answer the question below. Carefully solve the problem step by step, while using *as few words are possible*. Be careful about your choice of words. Use only necessary and essential steps. Avoid extra words. Avoid repetition. Avoid verbose statements. Avoid introductory remarks. When you’re ready to answer, please use the format ‘The answer is’.
F.1.3 User Input:
Question: [question here]
Solution:
F.2 Few-Shot Prompts
We provide 8-shot examples for each question, following the approach of Wei et al. (2022b), without any system instructions given to the language model. Each example is formatted in a chat-based template that includes both user and assistant roles. The order of examples for each question is randomly shuffled. We present one representative example for each prompt type, while the full set of examples is available in our public repository. Table 14-18 display one of eight examples used in FS-Human, FS-GPT4o, and FS-Self.
Total Replaced New
Method Before After Before After Ratio (%) After Ratio (%)
FS-Human 58.68 61.13 85.65 71.67 23.24 103.86 19.31
FS-GPT4o 67.06 66.26 82.39 67.76 35.25 113.07 17.61
FS-Self 71.88 68.98 81.52 67.72 43.14 109.34 15.34
FS-GPT4o-BoN 50.69 51.63 106.87 94.80 10.86 129.90 4.33
Table 7: Sample length statistics of selected shortest paths used for fine-tuning, before and after augmentation with zero-shot BoN samples, averaged across GSM8K and MATH and five main models. ‘Before’ and ‘After’ columns detail the average relative length of training data pre- and post-augmentation. ‘Ratio (%)’ indicates the proportion of each sample type after augmentation. Replaced refers to solutions for the same questions that were substituted based on shortest length selection. New denotes newly added samples from BoN that previously lacked correct solutions for each method.
Avg. Relative Ll-3B Ge-2B Qw-3B QM-1.5B DM-7B
Method Acc Len Acc Len Acc Len Acc Len Acc Len Acc Len
Zero-Shot Prompting
      Baseline 100.00 100.00 77.26 220.00 65.43 200.09 84.38 301.05 82.79 305.91 80.44 182.29
      Be Concise 99.85 88.46 78.24 168.35 64.22 163.14 83.78 258.16 82.94 304.14 80.74 180.58
      Hand Crafted 2 98.27 77.10 79.15 143.58 62.62 144.94 78.77 146.42 82.49 303.50 80.59 182.21
FT - External Data
      Direct Answer 24.88 1.36 17.36 2.12 11.30 3.42 22.90 3.43 22.97 3.46 23.96 3.42
      Human CoT 83.82 54.95 65.73 102.01 47.99 130.58 69.83 136.65 69.67 137.21 75.44 132.81
      GPT4o CoT 97.65 67.60 78.17 141.88 60.27 156.94 80.97 156.76 82.34 160.79 80.06 164.84
FT - Best-of-N Self-Generation
      Naive BoN 98.79 87.17 77.94 182.08 63.91 179.22 81.96 287.58 81.20 287.06 80.59 135.15
      Metareasoning 97.21 84.93 79.00 185.98 57.24 158.45 82.34 261.19 81.35 282.86 80.82 148.96
FS - Few-Shot Prompting
      FS-Human 96.18 64.37 77.71 115.14 59.74 155.41 78.09 233.80 79.00 111.19 81.27 141.84
      FS-GPT4o 97.52 77.88 77.03 146.81 60.35 168.01 79.15 296.32 83.02 163.29 81.73 158.39
      FS-Self 96.49 77.42 74.98 144.65 61.26 142.22 80.44 293.61 80.14 232.30 80.14 140.00
FT - Few-Shot Conditioned Self-Generation (ours)
      FS-Human 98.06 67.96 78.85 131.68 60.96 164.60 82.41 249.17 80.89 132.78 80.21 130.37
[…]
      FS-GPT4o 99.94 73.15 78.92 150.51 63.76 167.37 84.31 263.97 82.56 163.85 80.82 132.02
[…]
      FS-Self 98.86 77.51 78.47 142.74 62.09 159.44 82.94 270.71 82.34 243.01 80.52 134.24
[…]
FT - Few-Shot Conditioned Best-of-N Self-Generation (ours)
      FS-GPT4o-BoN 97.00 64.25 77.79 119.89 59.29 152.92 80.82 245.40 80.74 124.56 80.74 124.14
[…]
Table 8: Full results for main methods on GSM8K. Relative accuracy (%) and length (%) compared to default prompting are shown, averaged over five main models: Llama-3.2-3B, Gemma-2-2B, Qwen2.5-3B, Qwen2.5-Math-1.5B, and DeepSeekMath-7B. Absolute accuracy (%) and length (tokens) are reported for each individual model.
Avg. Relative Ll-3B Ge-2B Qw-3B QM-1.5B DM-7B
Method Acc Len Acc Len Acc Len Acc Len Acc Len Acc Len
Zero-Shot Prompting
      Baseline 100.00 100.00 44.40 511.75 24.00 375.99 56.80 605.45 69.80 551.36 37.00 357.32
      Be Concise 102.71 92.66 47.20 477.47 22.20 307.06 60.20 541.92 68.00 550.23 41.20 353.77
      Hand Crafted 2 101.62 85.26 46.60 445.85 24.00 275.69 54.80 393.80 68.80 543.72 40.00 365.17
FT - External Data
      Direct Answer 35.16 1.44 11.80 5.85 11.40 4.46 17.80 4.86 17.80 14.81 16.60 4.91
      Human CoT 75.61 53.14 26.80 233.95 21.20 240.38 38.20 218.88 49.40 273.08 33.80 251.39
      GPT4o CoT 90.52 87.21 35.80 381.65 26.00 410.97 56.60 395.13 49.20 407.29 34.60 403.96
FT - Best-of-N Self-Generation
      Naive BoN 101.74 89.89 47.20 433.42 21.60 341.47 60.00 559.09 70.00 521.26 39.40 311.08
      Metareasoning 103.02 90.56 46.20 421.56 23.60 348.31 56.20 539.82 69.60 508.67 42.20 344.43
FS - Few-Shot Prompting
      FS-Human 93.70 81.35 30.00 349.32 21.80 317.33 62.40 531.56 65.80 444.21 39.20 306.41
      FS-GPT4o 100.52 86.76 42.60 393.43 23.00 345.45 62.20 580.58 66.20 418.44 39.40 333.16
      FS-Self 103.28 85.30 46.20 407.70 20.60 323.20 64.40 572.48 71.80 473.90 40.80 287.19
FT - Few-Shot Conditioned Self-Generation (ours)
      FS-Human 99.69 87.78 42.00 411.29 22.20 357.87 59.80 538.92 68.00 502.23 40.20 297.41
[…]
      FS-GPT4o 101.87 87.58 45.20 397.24 22.80 351.49 59.20 560.48 69.20 495.32 40.40 301.53
[…]
      FS-Self 102.67 88.50 45.60 429.66 22.60 346.29 60.40 561.20 71.60 498.16 39.80 298.05
[…]
FT - Few-Shot Conditioned Best-of-N Self-Generation (ours)
      FS-GPT4o-BoN 102.56 76.30 43.60 316.12 24.00 313.52 60.60 523.00 67.40 378.64 41.20 290.39
[…]
Table 9: Full results for main methods on MATH. Relative accuracy (%) and length (%) compared to default prompting are shown, averaged over five main models: Llama-3.2-3B, Gemma-2-2B, Qwen2.5-3B, Qwen2.5-Math-1.5B, and DeepSeekMath-7B. Absolute accuracy (%) and length (tokens) are reported for each individual model.
Question
Cynthia eats one serving of ice cream every night. She buys cartons of ice cream with 15 servings of ice cream per carton at a cost of \$4.00 per carton. After 60 days, how much will she spend on ice cream?
Reasoning Paths
[Zero-Shot] To solve this problem, I’ll break it down into steps: Step 1: Determine the number of servings Cynthia eats in 60 days. Since Cynthia eats one serving of ice cream every night, she eats 1 serving per day. In 60 days, she will eat 60 servings. Step 2: Determine the number of cartons Cynthia needs to buy. Each carton has 15 servings of ice cream. To find out how many cartons Cynthia needs, I’ll divide the total number of servings she eats in 60 days (60) by the number of servings per carton (15). 60 / 15 = 4 So, Cynthia needs 4 cartons of ice cream. Step 3: Calculate the total cost of the ice cream. Each carton costs \$4.00. To find the total cost, I’ll multiply the number of cartons Cynthia needs (4) by the cost per carton (\$4.00). 4 * \$4.00 = \$16.00. Therefore, Cynthia will spend \$16.00 on ice cream after 60 days.
[FS-GPT4o-BoN] Cynthia eats one serving of ice cream every night, so in 60 days, she will eat 60 servings. Since each carton has 15 servings, she will need 60 / 15 = 4 cartons. Each carton costs \$4, so the total cost will be 4 cartons x \$4 per carton = \$16. The answer is \$16.
Table 10: Example rationales generated from the original Llama-3.2-3B model and the model fine-tuned using FS-GPT4o-BoN. The question is randomly sampled from GSM8k.
Question
Evaluate \$(1+2i)6-3i\$
Reasoning Paths
[Zero-Shot] Step 1: First, we need to distribute the 6 to both the 1 and the 2i in the expression. This means we multiply 6 by 1 and 6 by 2i. Step 2: Multiplying 6 by 1 gives us 6, and multiplying 6 by 2i gives us 12i. So, the expression becomes $6+12i-3i$. Step 3: Now, we need to combine like terms in the expression. The like terms in this expression are the real parts (6 and -3i) and the imaginary parts (12i and -3i). Step 4: Combining the real parts gives us 6 - 3i, and combining the imaginary parts gives us 12i - 3i = 9i. So, the expression simplifies to $6+9i$. The final answer is: $\boxed{6+9i}$
[FS-GPT4o-BoN] To evaluate the expression, we need to follow the order of operations (PEMDAS): 1. Multiply 6 by the complex number 1+2i: [ (1+2i)6 = 6 + 12i ] 2. Subtract 3i from the result: [ (6 + 12i) - 3i = 6 + 9i ]. The answer is 6 + 9i
Table 11: Example rationales generated from the original Llama-3.2-3B model and the model fine-tuned using FS-GPT4o-BoN. The question is randomly sampled from MATH500.
Question
While working at the restaurant, each of the forty customers who came into the restaurant gave Rafaela a \$20 tip. Julieta received 10% less money in tips than Rafaela. How much money did Julieta and Rafaela receive as tips altogether?
Reasoning Paths
[…]
[FS-GPT4o-BoN] Rafaela received \$20 x 40 = \$800 in tips. Julieta received 10 / 100 x \$800 = \$80 less than Rafaela. So, Julieta received \$800 - \$80 = \$720 in tips. Altogether, Rafaela and Julieta received \$800 + \$720 = \$1520 in tips. The answer is \$1520.
Table 12: Example rationales generated from the original Deekseek-7B model and the model fine-tuned using FS-GPT4o-BoN. The question is randomly sampled from GSM8k.
Question
Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$
Reasoning Paths
[Zero-Shot] To convert from rectangular coordinates to polar coordinates, we use the formulas $r=\sqrt{x^{2}+y^{2}}$ and $\theta=\arctan\frac{y}{x}$. In this case, $x=0$ and $y=3$, so $r=\sqrt{0^{2}+3^{2}}=3$ and $\theta=\arctan\frac{3}{0}$. However, $\arctan\frac{3}{0}$ is undefined, because we cannot divide by zero. Instead, we use the fact that $\theta$ is the angle that the line segment from the origin to the point $(0,3)$ makes with the positive $x$-axis. Since this line segment is vertical and points upwards, $\theta$ must be $\frac{\pi}{2}$. Therefore, the polar coordinates of the point $(0,3)$ are $(3,\frac{\pi}{2})$. The answer is $\boxed{(3,\frac{\pi}{2})}$.
[FS-GPT4o-BoN] We have that $r=\sqrt{0^{2}+3^{2}}=3$ and that $\theta=\arctan\frac{3}{0}=\frac{\pi}{2}.$ Thus, the polar coordinates are $\boxed{\left(3,\frac{\pi}{2}\right)}.$
Table 13: Example rationales generated from the original Deekseek-7B model and the model fine-tuned using FS-GPT4o-BoN. The question is randomly sampled from MATH500.
Table 14: One of eight human-annotated examples of concise reasoning for solving GSM8K and MATH problems from Wei et al. (2022b)
Question: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today?
Solution: There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6.
Table 15: One of eight GPT-4o-generated examples of concise reasoning for GSM8K problems.
Question: Aaron pays his actuary membership fees each year. The membership fee increases yearly by \$10. If he pays \$80 in the first year, how much does his membership cost, in dollars, in the sixth year?
Solution: 1. Start with the first year fee: \$80.
2. The fee increases by \$10 each year.
3. Calculate the fee for the second year: \$80 + \$10 = \$90.
4. Calculate the fee for the third year: \$90 + \$10 = \$100.
5. Calculate the fee for the fourth year: \$100 + \$10 = \$100.
6. Calculate the fee for the fifth year: \$100 + \$10 = \$103.
7. Calculate the fee for the sixth year: \$103 + \$10 = \$106. The answer is \$106.
Table 16: One of eight GPT-4o-generated examples of concise reasoning for MATH problems.
Question: A convex polyhedron $P$ has $26$ vertices, $60$ edges, and $36$ faces, $24$ of which are triangular and $12$ of which are quadrilaterals. A space diagonal is a line segment connecting two non-adjacent vertices that do not belong to the same face. How many space diagonals does $P$ have?
Solution: To find the number of space diagonals in the polyhedron, we first calculate the total number of line segments connecting any two vertices. This is given by the combination formula for choosing 2 vertices out of 26, which is: Total line segments = C(26, 2) = 26 * 25 / 2 = 325. Next, we subtract the number of edges and face diagonals from this total to find the number of space diagonals. The polyhedron has 60 edges, so we subtract these: 325 - 60 = 265. Now, we need to subtract the diagonals that lie on the faces. For the 24 triangular faces, there are no diagonals since a triangle has no diagonals. For the 12 quadrilateral faces, each quadrilateral has 2 diagonals. Therefore, the total number of face diagonals is: 12 * 2 = 24. Subtract these face diagonals from the remaining line segments: 265 - 24 = 241. Thus, the number of space diagonals is 241. The answer is 241.
Table 17: One of eight self-generated examples of concise reasoning for GSM8K problems from Llama-3.2-3B.
Question: A pie shop charges \$5 for a slice of pie. They cut each whole pie into 4 slices. How much money will the pie shop make if they sell 9 pies?
Solution: Here’s the step-by-step reasoning: 1. The pie shop cuts each whole pie into 4 slices, so they can sell 4 slices per pie.
2. They sell 9 pies, so they can sell a total of 9 x 4 = 36 slices.
3. Each slice costs \$5, so they will make 36 x \$5 = \$180. The answer is \$180.
Table 18: One of eight self-generated examples of concise reasoning for MATH problems from Llama-3.2-3B.
Question: When a polynomial is divided by $182x^{2}-7x+$ 18, what are the possible degrees of the remainder? Enter all the possible values, separated by commas.
Solution: ## Step 1: Recall the division algorithm for polynomials. When a polynomial $f(x)$ is divided by another polynomial $g(x),$ the remainder must be of degree less than $g(x).$# Step 2: Determine the degree of the divisor. The degree of $182x^{2}-7x+$ 18 is $2$# Step 3: Determine the possible degrees of the remainder. Since the remainder must be of degree less than $2,$ the possible degrees of the remainder are 00 and $1$ The final answer is: $boxed\{0,1\}$